# A quantum machine learning classifier to search for new physics

### v4

@@ 2.5 The numerical result
In the numerical example, we use $N=32$ two dimensional points which are divided into four classes equally. The training set is generated using scikit-learn Pedregosa:2011ork , and $x$, $y$ coordinates are translated as zenith and azimuth angles of the Bloch sphere, $\theta=(x+1)\pi/2$, $\varphi=(y+1)\pi$. When training, we use ‘Adam’ optimizer Kingma:2014vow , and the initialization of the ansatz parameters is randomly sampled from the many-body localized distribution Park:2024rim , the training is done by using the QuEST toolkit Jones:2019knd . When $l=3$, the ansatz is trained to a fidelity of $99.4\%$. In terms of quantum control instruction set (QCIS), the circuit for classification is depicted in Fig. 5. Then we scan the Bloch sphere and use QSN and VQSN to determine the classification assignment of each point on the Bloch sphere. The discrimination circuit is also run on tianyan176-2 of Tianyan quantum computing cloud platform, with Zuchongzhi superconducting quantum computing system Zhu:2021gkn .
@@ 3.1 Preparation of the datasets
The datasets are generated using MadGraph5 toolkitAlwall:2014hca ; Christensen:2008py ; Degrande:2011ua . A parton shower is applied using Pythia8 with NNPDF Sjostrand:2014zea ; Ball:2013hta , the fast detector simulation is applied using Delphes deFavereau:2013fsa with either the CMS or muon collider card, depending on the processes under study. The preparation of the datasets are applied using MLAnalysis Guo:2023nfu .
After the collision events are generated, a set of observables are picked to form the feature space, an event can be mapped into this feature space as a vector $\vec{v}$. Assuming $D$ observables are picked, then $\vec{v}$ is a $D$-dimensional vector. The z-score standardization Donoho_2004 is applied, $x^{j}_{m}=\left(v^{j}_{m}-\bar{v^{j}}\right)/\epsilon^{j}$, where $v^{j}_{m}$ is the $j$-th component ($0\leq j\leq D-1$) of $\vec{v}_{m}$ for the $m$-th collision event, $\bar{v^{j}}$ and $\epsilon^{j}$ are mean values and standard deviations of the $j$-th component of $\vec{v}$ over the training dataset. $\vec{x}$ is $\vec{v}$ after z-score standardization.
In the search for NP, the data to be dealt with are the collision events. The idea of this work is to decide whether some unknown events are neighbors of the SM events, or the NP events. In the case of HEP, there are different methods to define the distance between two events, which can have an impact on the results of event selection strategies Komiske:2019fks . In (V)QSN, we use the overlap of two quantum states to calculate the distance, therefore the definition of the distance depends on the method to encode vector $\vec{x}$ into a Hilbert space. There exist various methods for mapping vectors into a Hilbert space. Among these, the most straightforward approach is amplitude encoding. While different schemes may impact the efficiency of event selection, we employ the most straightforward amplitude encoding method for brevity and to maintain focus on (V)QSN, leaving the optimization of the encoding scheme for future work. For example, this study has not addressed improvements such as encoding the vectors in the dataset into a quantum state using a nonlinear approach Havlicek:2018nqz . It is also possible to utilize the vast Hilbert space of quantum computers to separate the signal from the background more when mapping to the quantum feature space Lloyd:2020eeh . The vector $\vec{x}$ is map to the quantum state,
@@ 3.2 Event selection strategy
The number of $|0\rangle_{b_{1}}$ outcome is denoted as $a$ and $\hat{a}=a/(rc)$ is used as a score to indicate the probability such how likely $\phi$ is from the NP, where $r\times c$ is the number of times $\langle\phi|\Psi\rangle$ is successfully built. And after $\hat{a}$ is obtained, we select events with $\hat{a}>\hat{a}_{th}$ where $\hat{a}_{th}$ is a tunable threshold that maximizes the signal significance which is defined as ParticleDataGroup:2020ssz ; Cowan:2010js ,
@@
where $N_{bg}$ and $N_{s}$ are number of background and signal events, respectively, and $N_{s,bg}=L\sigma_{s,bg}$, where $L$ is the luminosity, $\sigma_{s,bg}$ are the cross-sections of signal and background contributions. The accuracy of $\hat{a}$ can be estimated as $\mathcal{O}(1/\sqrt{rc})$.
For comparison with a classical algorithm, we also introduce the classical KNN algorithm which is already used in the search for NP DeSimone:2018efk . We use a modified KNN (denoted as ‘t-KNN’ where ‘t’ stands for tunable) with the following strategy for event selection.
@@ 3.3.1 The contribution of the gluon quartic gauge coupling at a muon collider
The gluon quartic gauge couplings (gQGCs) emerge from the Born-Infeld (BI) extension of the SM, which was originally driven by the idea that there should be an upper limit on the strength of the electromagnetic field Born:1934gh . It has been demonstrated that the BI model is also associated with M-theory-inspired models Ellis:2018cos ; Ellis:2021dfa ; Fradkin:1985qd ; Tseytlin:1999dj ; Cheung:2018oki . In the SMEFT, the operators contributing to gQGCs appear at dimension-8. As the importance of the dimension-8 operators in the SMEFT has been realized Green:2016trm ; Zhang:2020jyn ; Murphy:2020rsh ; Li:2020gnx ; Anders:2018oin ; Henning:2015alf , more and more phenomenological studies have been devoted to the dimension-8 operators in recent years Zhang:2018shp ; Bi:2019phv ; ATLAS:2017vqm ; CMS:2017rin ; CMS:2020ioi ; CMS:2016gct ; CMS:2017zmo ; CMS:2018ccg ; ATLAS:2018mxa ; CMS:2019uys ; CMS:2016rtz ; CMS:2017fhs ; Ellis:2023ucy ; Spor:2022hhn ; Spor:2022zob ; Yilmaz:2021ule ; Ellis:2020ljj ; Senol:2019swu ; Yilmaz:2019cue ; Ellis:2019zex ; Jahedi:2022duc ; Jahedi:2023myu . The case of gQGC is among the operators of dimension-8 in the SMEFT that have received attention. We focus on the $O_{gT,0}$ operator,
@@
where $G_{\mu\nu}^{a}$} is gluon field strengths, $W_{\mu\nu}^{i}$} denote electroweak field strengths, and $M_{0}$ is the NP scale. The expected constraints is $M_{0}\geq 1040\;{\rm MeV}$ at the LHC with the center of mass (c.m.) energy $\sqrt{s}=13\;{\rm TeV}$, and luminosity $L=36.7\;{\rm fb}^{-1}$ obtained by using the process $gg\to\gamma\gamma$ Ellis:2018cos . The combined sensitivities of the $Z\gamma$ and $\gamma\gamma$ channels at the LHC with $\sqrt{s}=13\;{\rm TeV}$ and $L=137\;{\rm fb}^{-1}$ are about three times of the above Ellis:2021dfa . For convenience, we define $f_{0}\equiv 1/(16M_{0}^{4})$.
At muon colliders, the gQGCs contributes to the process $\mu^{+}\mu^{-}\to\nu\bar{\nu}jj$ Yang:2023gos . The Feynman diagrams are shown in Fig. 7. This process is chosen as one of the processes to validate the VQSN since it is convenient to handle due to the absence of interference between the SM and NP.
@@ 3.3.2 Training of the ansatz
$2^{17}+2^{17}$ events are selected from the SM and NP events to form the training set. Since the $XY2M(\theta)=R_{z}(\theta)R_{x}(-\pi/2)R_{z}(-\theta)$ and $XY2P(\theta)=R_{z}(\theta)R_{x}(\pi/2)R_{z}(-\theta)$ are supported gates on tianyan176-2, the single-qubit layer of the ansatz are chosen to be $XY2M(\theta_{1})XY2P(\theta_{2})$ on each qubit. The entanglement layer is chosen to be pairwise. The ansatz is chosen to have $l=3$ entanglement layers, with $N=2^{18}$ and $d=4$, $n_{q}=20$ qubits are used. The ansatz is trained using QuEST, the fidelity reaches $78.5\%$.
@@
The test dataset are $50000+50000$ events selected from the SM and NP contributions, which are all different from the training events. In the case of running on the tianyan176-2, a sub set consist of $500+500$ events are selected. The circuit for discrimination in terms of QCIS on tianyan176-2 are shown in Fig. 8.
@@ 3.3.3 Numerical results
The receiver operating characteristic curve (ROC) represents a crucial metric for evaluating the efficacy of an classification algorithm. ROC is obtained by mapping the $\hat{a}_{th}$ ($\hat{b}_{th}$) into a two-dimensional plane with false positive rate (FPR) and true positive rate (TPR) as axes. The area under the ROC (AUC) can be utilized as a quantitative metric to assess the efficacy of a classification algorithm. The ROCs for (V)QSN and t-KNN are shown in Fig. 10. The large AUC in the case of tianyan176-2 can be attributed to the fact that the difference between all cases is small, therefore the results are affected by a limited number of events, and the test set running on tianyan176-2 consists of only $1\%$ events.
@@
where $\sigma^{\mu^{+}\mu^{-}\to jj\nu\bar{\nu}}_{\rm SM}$ and $\sigma^{\mu^{+}\mu^{-}\to jj\nu\bar{\nu}}_{\rm NP}$ are the cross-section of the SM and NP after requiring at least two jets, $\sigma^{\mu^{+}\mu^{-}\to jj\nu\bar{\nu}}_{\rm NP}$ is obtained when $f_{0}=\tilde{f}_{0}$. At $\sqrt{s}=3\;{\rm TeV}$ and $\tilde{f}_{0}=1\;{\rm TeV}^{-4}$, $\sigma^{\mu^{+}\mu^{-}\to jj\nu\bar{\nu}}_{\rm SM}=722.9\;{\rm fb}$ and $\sigma^{\mu^{+}\mu^{-}\to jj\nu\bar{\nu}}_{\rm NP}=3.21\;{\rm fb}$ Yang:2023gos . By choosing proper $\hat{a}_{th}$ and $\hat{b}_{th}$, the expected constraints on $|f_{0}|$ at $\sqrt{s}=3\;{\rm TeV}$ and $L=1\;{\rm ab}^{-1}$ (where $L$ is the luminosity chosen according to the design of the muon colliders Black:2022cth ; Accettura:2023ked ) are listed in Table 2. The expected constraints on $|f_{0}|$ by using a traditional event selection strategy at $\mathcal{S}_{start}=2$, $3$ and $5$ are <0.155<0.155, <0.191<0.191 and $0.248\;{\rm TeV}^{-4}$, respectively. The dataset in this paper is in fact a subset (about $28\%$) of the one used in Ref. Yang:2023gos , it can be seen that the VQSN can reach a same order of magnitude, while t-KNN outperforms the traditional method.
@@
The cc of VQSN and its classical counterpart, i.e. (t-)KNN is compared. For one test event, $168$ gates need to be applied in terms of QCIS, therefore, when $c=1000$, the cc is estimated as $168000$. For (t-)KNN, the cc is for one test event is $1.57\times 10^{6}$ because the training set contains $2^{18}$ 66-dimensional vectors. It can be seen that, even with a classical dataset, the cc of VQSN can be one order of magnitude smaller.
@@ 3.4.1 The gluon quartic gauge couplings and the process to be considered
One of the reasons muon collider is better suited to the search for NP is that the QCD background is suppressed. ML algorithms are more suitable for processes with more intricate final states, which necessitates greater investment in kinematic analysis when traditional methods are employed. There are certain complexities involved in discussing QCD background at the LHC, which offer an opportunity for utilizing ML algorithms in this regard. Therefore, we also investigate an example of the search search for gQGCs at the LHC.
@@
Except for the SM contribution, other processes contributing to background due to detector effects are considered, including $pp\to(\gamma)\ell\ell+jets$ with missing jets, $pp\to t\bar{t}\to jjW^{+}W^{-}\to jj\ell^{+}\ell^{-}\nu\bar{\nu}$ with b-jets mis-tagged, and $pp\to j\ell\ell\ell\nu$ with a missing charged lepton. They are listed in Table 3, with the cross-sections obtained by using MadGraph5 toolkit Alwall:2014hca , except for the case of $t\bar{t}+jets$, in which we use inclusive cross-section of $t\bar{t}+jets$ production Czakon:2011xx ; CMS:2020grm ; ATLAS:2024aht .
@@ 3.4.2 Training of the ansatz
The training dateset is consist of events after $N_{part}$ cut. The number of events picked to form the training dataset is listed in Table 5. Training dataset has $N=65536$ vectors, with three qubits to encode one vector, with $d=8$, the number of qubits to store $\Psi_{\rm train}$ (or the ansatz $\Psi(\alpha_{i})$) is $n_{q}=19$. In this subsection, the ‘shifted-circular-alternating’ (SCA) entanglement layer Sim:2019yyv is used. The circuit of the ansatz is the one shown in Fig. 4.

### v5

@@ 2.5 The numerical result
In the numerical example, we use $N=32$ two dimensional points which are divided into four classes equally. The training set is generated using scikit-learn Pedregosa and others (2011), and $x$, $y$ coordinates are translated as zenith and azimuth angles of the Bloch sphere, $\theta=(x+1)\pi/2$, $\varphi=(y+1)\pi$. When training, we use ‘Adam’ optimizer Kingma and Ba (2014), and the initialization of the ansatz parameters is randomly sampled from the many-body localized distribution Park et al. (2024), the training is done by using the QuEST toolkit Jones et al. (2019). When $l=3$, the ansatz is trained to a fidelity of $99.4\%$. In terms of quantum control instruction set (QCIS), the circuit for classification is depicted in Fig. 5. Then we scan the Bloch sphere and use QSN and VQSN to determine the classification assignment of each point on the Bloch sphere. The discrimination circuit is also run on tianyan176-2 of Tianyan quantum computing cloud platform, with Zuchongzhi superconducting quantum computing system Zhu and others (2022).
@@ 3.1 Preparation of the datasets
The datasets are generated using MadGraph5 toolkitAlwall et al. (2014); Christensen and Duhr (2009); Degrande et al. (2012). A parton shower is applied using Pythia8 with NNPDF Sjöstrand et al. (2015); Ball et al. (2013), the fast detector simulation is applied using Delphes de Favereau et al. (2014) with either the CMS or muon collider card, depending on the processes under study. The preparation of the datasets are applied using MLAnalysis Guo et al. (2024).
After the collision events are generated, a set of observables are picked to form the feature space, an event can be mapped into this feature space as a vector $\vec{v}$. Assuming $D$ observables are picked, then $\vec{v}$ is a $D$-dimensional vector. The z-score standardization Donoho and Jin (2004) is applied, $x^{j}_{m}=\left(v^{j}_{m}-\bar{v^{j}}\right)/\epsilon^{j}$, where $v^{j}_{m}$ is the $j$-th component ($0\leq j\leq D-1$) of $\vec{v}_{m}$ for the $m$-th collision event, $\bar{v^{j}}$ and $\epsilon^{j}$ are mean values and standard deviations of the $j$-th component of $\vec{v}$ over the training dataset. $\vec{x}$ is $\vec{v}$ after z-score standardization.
In the search for NP, the data to be dealt with are the collision events. The idea of this work is to decide whether some unknown events are neighbors of the SM events, or the NP events. In the case of HEP, there are different methods to define the distance between two events, which can have an impact on the results of event selection strategies Komiske et al. (2019). In (V)QSN, we use the overlap of two quantum states to calculate the distance, therefore the definition of the distance depends on the method to encode vector $\vec{x}$ into a Hilbert space. There exist various methods for mapping vectors into a Hilbert space. Among these, the most straightforward approach is amplitude encoding. While different schemes may impact the efficiency of event selection, we employ the most straightforward amplitude encoding method for brevity and to maintain focus on (V)QSN, leaving the optimization of the encoding scheme for future work. For example, this study has not addressed improvements such as encoding the vectors in the dataset into a quantum state using a nonlinear approach Havlicek et al. (2019). It is also possible to utilize the vast Hilbert space of quantum computers to separate the signal from the background more when mapping to the quantum feature space Lloyd et al. (2020). The vector $\vec{x}$ is mapped to the quantum state,
@@ 3.2 Event selection strategy
The number of $|0\rangle_{b_{1}}$ outcome is denoted as $a$ and $\hat{a}=a/(rc)$ is used as a score to indicate the probability such how likely $\phi$ is from the NP, where $r\times c$ is the number of times $\langle\phi|\Psi\rangle$ is successfully built. And after $\hat{a}$ is obtained, we select events with $\hat{a}>\hat{a}_{th}$ where $\hat{a}_{th}$ is a tunable threshold that maximizes the signal significance which is defined as Zyla and others (2020); Cowan et al. (2011),
@@
where $N_{bg}$ and $N_{s}$ are numbers of background and signal events, respectively, and $N_{s,bg}=L\sigma_{s,bg}$, where $L$ is the luminosity, $\sigma_{s,bg}$ are the cross-sections of signal and background contributions. The accuracy of $\hat{a}$ can be estimated as $\mathcal{O}(1/\sqrt{rc})$.
For comparison with a classical algorithm, we also introduce the classical KNN algorithm which is already used in the search for NP De Simone and Jacques (2019). We use a modified KNN (denoted as ‘t-KNN’ where ‘t’ stands for tunable) with the following strategy for event selection.
@@ 3.3.1 The contribution of the gluon quartic gauge coupling at a muon collider
The gluon quartic gauge couplings (gQGCs) emerge from the Born-Infeld (BI) extension of the SM, which was originally driven by the idea that there should be an upper limit on the strength of the electromagnetic field Born and Infeld (1934). It has been demonstrated that the BI model is also associated with M-theory-inspired models Ellis and Ge (2018); Ellis et al. (2022); Fradkin and Tseytlin (1985); Tseytlin (1999); Cheung et al. (2018). In the SMEFT, the operators contributing to gQGCs appear at dimension-8. As the importance of the dimension-8 operators in the SMEFT has been realized Green et al. (2017); Zhang and Zhou (2020); Murphy (2020); Li et al. (2021); Anders and others (2018); Henning et al. (2017), more and more phenomenological studies have been devoted to the dimension-8 operators in recent years Zhang and Zhou (2019); Bi et al. (2019); Aaboud and others (2017); Khachatryan and others (2017b); Sirunyan and others (2020a); Khachatryan and others (2017a); Sirunyan and others (2017, 2019c); Aaboud and others (2019); Sirunyan and others (2019a); Khachatryan and others (2016); Sirunyan and others (2018); Ellis et al. (2023); Spor (2023); Spor et al. (2022); Yilmaz (2021); Ellis et al. (2021); Senol et al. (2019); Yilmaz et al. (2020); Ellis et al. (2020); Jahedi and Lahiri (2023); Jahedi (2023). The case of gQGC is among the operators of dimension-8 in the SMEFT that have received attention. We focus on the $O_{gT,0}$ operator,
@@
where $G_{\mu\nu}^{a}$} is gluon field strengths, $W_{\mu\nu}^{i}$} denote electroweak field strengths, and $M_{0}$ is the NP scale. The expected constraint is $M_{0}\geq 1040\;{\rm MeV}$ at the LHC with the center of mass (c.m.) energy $\sqrt{s}=13\;{\rm TeV}$, and luminosity $L=36.7\;{\rm fb}^{-1}$ obtained by using the process $gg\to\gamma\gamma$ Ellis and Ge (2018). The combined sensitivities of the $Z\gamma$ and $\gamma\gamma$ channels at the LHC with $\sqrt{s}=13\;{\rm TeV}$ and $L=137\;{\rm fb}^{-1}$ are about three times of the above Ellis et al. (2022). For convenience, we define $f_{0}\equiv 1/(16M_{0}^{4})$.
At muon colliders, the gQGCs contribute to the process $\mu^{+}\mu^{-}\to\nu\bar{\nu}jj$ Yang et al. (2023a). The Feynman diagrams are shown in Fig. 7. This process is chosen as one of the processes to validate the VQSN since it is convenient to handle due to the absence of interference between the SM and NP.
@@ 3.3.2 Training of the ansatz
$2^{17}+2^{17}$ events are selected from the SM and NP events to form the training set. Since the $XY2M(\theta)=R_{z}(\theta)R_{x}(-\pi/2)R_{z}(-\theta)$ and $XY2P(\theta)=R_{z}(\theta)R_{x}(\pi/2)R_{z}(-\theta)$ are supported gates on tianyan176-2, the single-qubit layers of the ansatz are chosen to be $XY2M(\theta_{1})XY2P(\theta_{2})$ on each qubit. The entanglement layer is chosen to be pairwise. The ansatz is chosen to have $l=3$ entanglement layers, with $N=2^{18}$ and $d=4$, $n_{q}=20$ qubits are used. The ansatz is trained using QuEST, the fidelity reaches $78.5\%$.
@@
The test dataset is $50000+50000$ events selected from the SM and NP contributions, which are all different from the training events. In the case of running on the tianyan176-2, a sub set consist of $500+500$ events are selected. The circuit for discrimination in terms of QCIS on tianyan176-2 is shown in Fig. 8.
@@ 3.3.3 Numerical results
The receiver operating characteristic curve (ROC) represents a crucial metric for evaluating the efficacy of a classification algorithm. ROC is obtained by mapping the $\hat{a}_{th}$ ($\hat{b}_{th}$) into a two-dimensional plane with false positive rate (FPR) and true positive rate (TPR) as axes. The area under the ROC (AUC) can be utilized as a quantitative metric to assess the efficacy of a classification algorithm. The ROCs for (V)QSN and t-KNN are shown in Fig. 10. The large AUC in the case of tianyan176-2 can be attributed to the fact that the difference between all cases is small, therefore the results are affected by a limited number of events, and the test set running on tianyan176-2 consists of only $1\%$ events.
@@
where $\sigma^{\mu^{+}\mu^{-}\to jj\nu\bar{\nu}}_{\rm SM}$ and $\sigma^{\mu^{+}\mu^{-}\to jj\nu\bar{\nu}}_{\rm NP}$ are the cross-section of the SM and NP after requiring at least two jets, $\sigma^{\mu^{+}\mu^{-}\to jj\nu\bar{\nu}}_{\rm NP}$ is obtained when $f_{0}=\tilde{f}_{0}$. At $\sqrt{s}=3\;{\rm TeV}$ and $\tilde{f}_{0}=1\;{\rm TeV}^{-4}$, $\sigma^{\mu^{+}\mu^{-}\to jj\nu\bar{\nu}}_{\rm SM}=722.9\;{\rm fb}$ and $\sigma^{\mu^{+}\mu^{-}\to jj\nu\bar{\nu}}_{\rm NP}=3.21\;{\rm fb}$ Yang et al. (2023a). By choosing proper $\hat{a}_{th}$ and $\hat{b}_{th}$, the expected constraints on $|f_{0}|$ at $\sqrt{s}=3\;{\rm TeV}$ and $L=1\;{\rm ab}^{-1}$ (where $L$ is the luminosity chosen according to the design of the muon colliders Black and others (2024); Accettura and others (2023)) are listed in Table 2. The expected constraints on $|f_{0}|$ by using a traditional event selection strategy at $\mathcal{S}_{start}=2$, $3$ and $5$ are <0.155<0.155, <0.191<0.191 and $0.248\;{\rm TeV}^{-4}$, respectively. The dataset in this paper is in fact a subset (about $28\%$) of the one used in Ref. Yang et al. (2023a), it can be seen that the VQSN can reach a same order of magnitude, while t-KNN outperforms the traditional method.
@@
The cc of VQSN and its classical counterpart, i.e. (t-)KNN are compared. For one test event, $168$ gates need to be applied in terms of QCIS, therefore, when $c=1000$, the cc is estimated as $168000$. For (t-)KNN, the cc is for one test event is $1.57\times 10^{6}$ because the training set contains $2^{18}$ 66-dimensional vectors. It can be seen that, even with a classical dataset, the cc of VQSN can be one order of magnitude smaller.
@@ 3.4.1 The gluon quartic gauge couplings and the process to be considered
One of the reasons muon collider is better suited to the search for NP is that the QCD background is suppressed. ML algorithms are more suitable for processes with more intricate final states, which necessitates greater investment in kinematic analysis when traditional methods are employed. There are certain complexities involved in discussing QCD background at the LHC, which offer an opportunity for utilizing ML algorithms in this regard. Therefore, we also investigate an example of the search for gQGCs at the LHC.
@@
Except for the SM contribution, other processes contributing to background due to detector effects are considered, including $pp\to(\gamma)\ell\ell+jets$ with missing jets, $pp\to t\bar{t}\to jjW^{+}W^{-}\to jj\ell^{+}\ell^{-}\nu\bar{\nu}$ with b-jets mis-tagged, and $pp\to j\ell\ell\ell\nu$ with a missing charged lepton. They are listed in Table 3, with the cross-sections obtained by using MadGraph5 toolkit Alwall et al. (2014), except for the case of $t\bar{t}+jets$, in which we use inclusive cross-section of $t\bar{t}+jets$ production Czakon and Mitov (2014); Sirunyan and others (2020b); Aad and others (2024).
@@ 3.4.2 Training of the ansatz
The training dateset consists of events after $N_{part}$ cut. The number of events picked to form the training dataset is listed in Table 5. Training dataset has $N=65536$ vectors, with three qubits to encode one vector, with $d=8$, the number of qubits to store $\Psi_{\rm train}$ (or the ansatz $\Psi(\alpha_{i})$) is $n_{q}=19$. In this subsection, the ‘shifted-circular-alternating’ (SCA) entanglement layer Sim et al. (2019) is used. The circuit of the ansatz is the one shown in Fig. 4.